# Disturbance Compensation for Safe Kinematic Control of Robotic Systems with Closed Architecture

### v1

@@ Abstract
In commercial robotic systems, it is common to encounter a closed inner-loop torque controller that is not user-modifiable. However, the outer-loop controller, which sends kinematic commands such as position or velocity for the inner-loop controller to track, is typically exposed to users. In this work, we focus on the development of an easily integrated add-on at the outer-loop layer by combining disturbance rejection control and robust control barrier function for high-performance tracking and safe control of the whole dynamic system of an industrial manipulator. This is particularly beneficial when 1) the inner-loop controller is imperfect, unmodifiable, and uncertain; and 2) the dynamic model exhibits significant uncertainty. Stability analysis, formal safety guarantee proof, and hardware experiments with a PUMA robotic manipulator are presented. Our solution demonstrates superior performance in terms of simplicity of implementation, robustness, tracking precision, and safety compared to the state of the art. Video: https://youtu.be/zw1tanvrV8Q
@@ I Introduction
Robotic systems typically use a hierarchical architecture combining perception, planning, and low-level control, which improves modularity and reliability. In the control layer, a kinematic controller (outer loop) and a dynamic controller (inner loop) are commonly used. Since tuning the inner loop requires expertise, it is usually hidden from users for safety, forming a closed architecture [1, 2, 3, 4]. Users can thus only design the kinematic controller, sending position or velocity commands. However, wear, modifications, or payload changes can alter dynamics and degrade inner-loop performance [5]. This work addresses the challenge of achieving unified high-performance and safe tracking control using only the kinematic controller, despite an imperfect, unmodifiable inner loop, and uncertain dynamics.
Fig. 1 shows that our approach successfully follows an infinity-shaped reference trajectory despite various uncertainties and a poor inner-loop unknown controller (see the blue trajectory). In contrast, using the poor inner-loop controller without our approach’s compensation results in poor trajectory tracking (see the red trajectory).
Figure 1: Trajectory tracking of the PUMA 500 for an infinity-shaped reference.
@@
A detailed literature review can be found in the subsequent section. Existing works typically rely on the following assumptions: 1) The inner-loop controller is known and accessible to the user; 2) The inner-loop controller is known but not accessible to the user; 3) The inner-loop controller has perfect tracking for outer-loop commands. Our goal is to minimize reliance on knowledge of the dynamic model and the inner-loop controller, while still guaranteeing tracking and safe control using only an outer-loop kinematic controller.
As shown in Fig. 2, existing robotic systems typically send desired position ($q^{*}$) or velocity ($\dot{q}^{*}$) commands to the inner-loop controller. Our solution, inserted between them (gray box), comprises a disturbance-rejection tracking controller, an extended state observer (ESO), and an extended state observer-based robust quadratic programming (ESOR-QP) module. The ESO estimates the total disturbance $\hat{f}$ using the joint angle $q$ and the safe kinematic control $q_{d,\text{safe}}$ or $\dot{q}_{d,\text{safe}}$. This estimate is compensated in the tracking controller to produce nominal commands $q_{d}$ or $\dot{q}_{d}$, which are monitored by the ESOR-QP to enforce robust safety constraints.
@@
  • We innovatively formulate the kinematic control of closed-architecture robotic manipulator systems as a disturbance rejection problem, requiring only a practical assumption on the structure, Jacobian, and nominal inertia matrix.
  • We develop a robust control barrier function (CBF) framework based on an extended state observer, enabling safe control of general nonlinear affine systems with high relative degrees under disturbances. To the best of our knowledge, this is the first work achieving safe kinematic control of closed-architecture robots under largely unknown uncertainties.
  • As shown in Fig. 2, our method is implemented as a standalone add-on module that enhances existing robotic manipulators without modifying the inner-loop controller. Real-time control at 1 kHz is demonstrated on a PUMA 500 manipulator.
@@
The rest of the paper is organized as follows: Section II reviews the related work. Section III introduces the preliminaries and problem formulation. Section IV presents the proposed framework. Hardware experiments are reported in Section V. Concluding remarks are in Section VI.
@@ II Related Work
An overview of the comparison between our work and the most related works is presented in Table I.
II-A Adaptive Control for Closed Architecture
Learning-based Adaptive Control: [2, 3, 4] employ neural network–based outer-loop adaptation to approximate the kinematics, dynamics, and inner-loop controller. As summarized in Table I, they make the weakest assumption about the inner loop, identical to ours, where $\Psi(\cdot)$ denotes an unknown structure possibly containing proportional, integral, or state-dependent terms. However, they rely on numerous tuning parameters and limited interpretability.
Conventional Adaptive Control: [1] introduces dynamic modularity with an adaptive controller under a moderate assumption of a known PI/PID inner-loop structure.
These studies address tracking control only, without safety considerations. In our experiments, we compare tracking performance with the learning-based adaptive control due to the shared inner-loop assumption.
@@ II-A Adaptive Control for Closed Architecture
[2, 3, 4]
@@
[6, 7, 8]
Robust CBF - - [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]
@@ II-B Reduced-order Model-based CBF
Many CBF-based safety approaches rely on high-fidelity dynamic models with torque control, yet obtaining such models remains challenging. Reduced-order or kinematic CBFs have been introduced to ensure safety with lower-fidelity models. For instance, [6] guarantees safety via safe velocity control but assumes a perfect inner-loop controller capable of instant tracking. Similarly, [7, 8] employ kinetic energy–based CBFs that offer partial robustness to model uncertainty but still depend on numerous dynamic parameters and a known PD-type inner loop. In contrast, our work addresses closed-architecture robots with minimal knowledge of the inner-loop controller and significant model uncertainty.
@@ II-C Robust CBF
Robust CBFs have recently been developed to handle model uncertainties and external disturbances [9, 10, 11, 12, 13, 14, 15, 17, 19, 20, 21, 25, 18, 16, 23, 24, 22]. Among them, disturbance observer–based robust CBFs (DOB-CBFs) [25, 18, 16, 23, 24, 22, 15, 17, 19, 20, 21] enable active disturbance estimation without requiring worst-case bounds. Unlike these methods where estimation error bounds are absent or loosely derived via Lyapunov analysis, our approach obtains them directly from the observer error dynamics, resulting in tighter and less conservative bounds. Experimental comparisons with state-of-the-art methods (e.g., [16]) further demonstrate this advantage. In contrast to our conference version [25] that only considered relative-degree-one systems, this work provides a generalized framework with rigorous proofs guaranteeing the forward invariance of the safety set.
@@ III-A System Dynamics
The most common inner-loop controller for robot manipulators typically consists of a PID controller or PD controller, often combined with dynamic compensation, such as gravity. For stability, most inner-loop controllers include a derivative term for damping, $(\dot{q}-\dot{q}_{d})$, where $\dot{q}_{d}$ is a joint velocity command.
@@
with $K_{d}$ a proportional control gain, and $\Psi(\cdot)$ is a general, unknown function that may include proportional, integral, derivative, and dynamic compensation terms. To the best of our knowledge, this assumption is the most general in the literature (see Table I), requiring minimal knowledge of the inner-loop controller, similar to [2], and more general than [1].
@@
The system dynamic model combining (1) and (2) is:
$\begin{split}\ddot{q}&={M}^{-1}\Bigl(-{C}\dot{q}-{G}-F_{r}+{J}^{T}F_{ext}-{K}_{d}\dot{q}+\Psi(\cdot)+{K}_{d}\dot{q}_{d}\Bigr)\end{split}.$ (3)
We have removed torque control, as it is not accessible in a closed architecture. $\dot{q}_{d}$ is the kinematic control to be manipulated in the outer-loop layer, and its control law will be designed. Another reason for choosing $\dot{q}_{d}$ instead of $q_{d}$ as the control input is to avoid differentiation-induced noise.
@@ III-B Robot Kinematic Control with Known Model
Neglecting $F_{r}$ and $F_{ext}$ in (1), the exact system dynamics is used to design the control law:
@@
where $\ddot{q}^{\star},\dot{q}^{\star},q^{\star}\in\mathbb{R}^{n}$ represent the desired acceleration, velocity, and position of a reference joint trajectory generated by an upstream planner. $k_{p}\in\mathbb{R}^{n\times n}$ and $k_{d}\in\mathbb{R}^{n\times n}$ are the proportional and derivative gain matrices, respectively. $\ddot{q}^{\star}$ serves only as a feedforward term, while $u_{0}$ is a position controller. By substituting (4) and (5) into (3), we get:
@@
Proper $k_{p}$ and $k_{d}$ make q{q} converge to ${q}^{\star}$ asymptotically.
@@ III-C Control Barrier Functions
The system (3) can be rewritten as a nonlinear control-affine system when ignoring the disturbance vector $F_{ext}$:
@@
The safety of system (7) can be guaranteed using a safety set. A set $\mathcal{C}$ is considered to be a safety set if it is forward invariant in the state space $\mathcal{X}$, i.e., for system (7) if solutions for some $u\in\mathcal{U}$ starting at any initial safe condition $x(0)\in\mathcal{C}$ satisfy $x(t)\in\mathcal{C}$, $\forall t\geq 0$. The safety set $\mathcal{C}$ is defined as a 0-superlevel set of a continuously-differential function $h(x):\mathbb{R}^{2n}\rightarrow\mathbb{R}$ as:
@@
(Exponential CBF (ECBF) [26, 27]) Consider system (7) with relative degree $r$ for an $r$-times continuously differentiable function $h$, i.e., $L_{g}L_{\psi}h(x)=\cdots=L_{g}L_{\psi}^{r-2}h(x)=0$ and $L_{g}L_{\psi}^{r-1}h(x)\neq 0,\forall x\in\mathcal{C}$. $h(x)$ is an ECBF if there exists a row vector $K_{a}\in\mathbb{R}^{r}$ satisfying $\forall x\in\mathcal{C}$
@@
where $\eta_{b}(x)=[h(x),\dot{h}(x),\cdots,h^{(r-1)}(x)]^{T}$, $K_{a}=[k_{1},\cdots,k_{r}]$, and the values of $k_{1},\cdots,k_{r}$ satisfy specific properties given in [26, 27].
To guarantee safety, the control problem is formulated as a quadratic program (QP) with a CBF as a hard constraint [28]:
@@
Research Objective: Given the system in (3), the nominal dynamics with limited knowledge of the system and inner-loop controller are defined as
$\ddot{q}=\bar{M}^{-1}\!\left(-\bar{C}\dot{q}-\bar{G}+\bar{K}_{d}\dot{q}_{d}\right),$ (11)
where $\bar{M}$, $\bar{C}$, and $\bar{G}$ are nominal parameters, and $\bar{K}_{d}$ is the nominal inner-loop gain. The goal is to design a $\dot{q}_{d}$ controller that achieves accurate trajectory tracking and task-space safety (e.g., collision avoidance) under an unknown inner-loop controller, uncertain dynamics, and external disturbances.
@@
We use the nominal model (11) for general mathematical derivation, whereas the reduced nominal model (12) is employed in hardware experiments to evaluate the tracking performance and safety under limited information.

### v2

@@ Abstract
In commercial robotic systems, it is common to encounter a closed inner-loop (low-level) torque controller that is not user-modifiable. However, the outer-loop controller, which sends kinematic commands such as position or velocity for the inner-loop controller to track, is typically exposed to users. In this work, we focus on the development of an easily integrated add-on at the outer-loop layer by combining disturbance rejection control and robust control barrier function for high-performance tracking and safe control of the whole dynamic system of an industrial manipulator. This is particularly beneficial when 1) the inner-loop controller is imperfect, unmodifiable, and uncertain; and 2) the dynamic model exhibits significant uncertainty. Stability analysis, formal safety guarantee proof, simulations, and hardware experiments with a PUMA robotic manipulator are presented. Our solution demonstrates superior performance in terms of simplicity of implementation, robustness, tracking precision, and safety compared to the state of the art. Video: https://youtu.be/zw1tanvrV8Q
@@ I Introduction
Robotic systems often employ hierarchical software design, stacking perception, decision-making, planning, and low-level control. Such modularity is particularly beneficial for troubleshooting and improving the reliability of robotic systems. For example, in the control block, a combination of a kinematic controller (outer-loop controller) and a dynamic controller (inner-loop controller) is commonly seen in various robots. However, because tuning the inner-loop controller requires expert knowledge, this component is typically not exposed to users due to product safety considerations, a practice referred to as closed architecture in the literature [wang2019dynamic, ahanda2022adaptive, ahanda2022task, khan2024control]. In other words, users are only allowed to design the kinematic controller, sending position or velocity for the inner-loop controller to track. Additionally, mechanical parts can wear out over time, and users may modify the robot by adding or replacing parts or attaching payloads, which can lead to changes in dynamics and potentially render the inner-loop controller suboptimal [liu2022advancements]. The fundamental challenge we address in this article is how to achieve high-performance tracking control and safe control in a unified framework for the entire dynamic system using only the kinematic controller, while the inner-loop controller is imperfect and unmodifiable, and both the dynamic model and the inner-loop controller exhibit significant uncertainty. Fig. 1 shows that our approach successfully follows an infinity-shaped reference trajectory despite various uncertainties and a poor inner-loop unknown controller (see the blue trajectory). In contrast, using the poor inner-loop controller without our approach’s compensation results in poor trajectory tracking (see the red trajectory).
Figure 1: Trajectory tracking of the PUMA 500 for an infinity-shaped reference. The trajectory using a poor inner-loop controller is shown in red, while the improved tracking performance (ours) is shown in blue.
To bear in mind the two control objectives: high-performance tracking and safe control, we briefly survey the literature in the categories of adaptive control and reduced-order model-based barrier functions to illustrate our current position. A detailed literature review can be found in the subsequent section. [wang2019dynamic] presents a class of adaptive outer-loop control schemes that generate joint velocity and position commands for the inner-loop controller with certain proportional-integral (PI) or proportional-integral-derivative (PID) structures. [ahanda2022adaptive, khan2024control] relaxes the inner-loop controller’s structure and employs neural networks-based adaptive control to approximate the inner-loop controller’s dynamics and both kinematic and dynamic parameters. These works only consider tracking control without safety functions, such as state constraints. Additional limitations related to assumptions, implementation complexity, and performance will be discussed in the literature review and experimental comparisons.
Control barrier functions (CBFs) have emerged as a powerful tool for designing controllers that guarantee safety [ames2016control, ames2019control]. However, most existing CBF approaches rely on high-fidelity dynamic models. Reduced-order model-based CBFs [molnar2021model] demonstrate robust obstacle avoidance using only safe velocities, without requiring high-fidelity dynamic models. However, [molnar2021model] assumes perfect inner-loop tracking performance. [singletary2021safety] and [molnar2023safety] propose CBFs with kinetic energy and input-to-state safe backup set methods, respectively, reducing model dependence and allowing for some model uncertainty. However, they still rely on prior knowledge of numerous dynamic model parameters and a known-structure inner-loop controller.
@@
To sum up, existing works typically rely on the following assumptions: 1) The inner-loop controller is known and accessible to the user; 2) The inner-loop controller is known but not accessible to the user; 3) The inner-loop controller has perfect tracking for outer-loop commands. Our goal is to minimize reliance on knowledge of the dynamic model and the inner-loop controller, while still guaranteeing tracking and safe control using only an outer-loop kinematic controller.
We develop a novel add-on at the kinematic control layer, by combining extended state observer (ESO)-based disturbance rejection control and robust CBF, called ESOR-QP [chen2023robust] for high-performance tracking and provably safe control. Distinct from the aforementioned methods, our method treats the imperfect inner-loop controller, unmodeled dynamics, and external disturbance as a total disturbance, estimated by ESO, and then sends it to both the disturbance rejection controller and the robust CBF. Note that our solution is not simply a merge of two functionalities in tracking and safe control. Instead, a well-estimated total disturbance is naturally used as compensation for disturbance rejection and safe control. The so-called compensation has dual benefits in these two components. For trajectory tracking in task space, the tracking controller rejects the total disturbance for high-performance trajectory tracking. For safe control to avoid collisions, the estimated disturbance is integrated into a robust constraint in the CBF-based optimal control.
As shown in Fig. 2, our developed outer-loop controller is composed of a disturbance rejection tracking controller, an ESOR-QP module, and an ESO module. First, the ESO uses output measurement $q$ (joint angle) and final safe kinematic control $q_{d_{safe}}$ (position control) or $\dot{q}_{d_{safe}}$ (velocity control) to estimate the total disturbance $\hat{f}$. Second, the estimated total disturbance $\hat{f}$ is rejected in the tracking controller to generate outer-loop kinematic control commands $q_{d}$ or $\dot{q}_{d}$. Third, our robust CBF, ESOR-QP, continuously monitors the nominal control $q_{d}$ or $\dot{q}_{d}$ and intervenes if the nominal control is deemed unsafe. Such a safety filter relies on robust safety constraints integrating the nominal model as well as the estimated disturbance $\hat{f}$.
@@
  • System-level contribution: We propose a unified control framework that enhances tracking performance and safety using only a kinematic controller. Our approach does not require access to the inner-loop controller and remains effective in the presence of model parameter uncertainties, a largely unknown inner-loop controller, and external disturbances from the environment.
  • Theoretical contribution: We develop a robust CBF based on an extended state observer, enabling safe control of general nonlinear affine systems with high-order relative degrees under disturbances.
  • Practical contribution: As illustrated in Fig. 2, our solution is implemented as a standalone add-on component that can be integrated into existing legacy robotic manipulators to improve performance without modifying the inner-loop controller. Real-time control at 1 kHz has been implemented on a PUMA 500 robotic manipulator.
@@
We have made the following substantial theoretical and experimental improvements over our preliminary conference version [chen2023robust]:
  1. 1.
    Theory: We have extended our robust CBF for a system with relative degree one to accommodate arbitrary relative degrees, i.e., higher-order robust CBFs. Our work provides rigorous proof, complementing [chen2023robust], for the forward invariance property of the safety set.
  2. 2.
    Application: We demonstrate the efficacy of our approach through validation in real-world hardware experiments on a nontrivial robotic manipulator, addressing the particular challenge of a closed architecture. In contrast, our previous work was validated on simpler robotic systems in simulation, including a cruise control system with 1D control and a self-balancing scooter with 2D control.
The rest of the paper is organized as follows: Section II reviews the related work. Section III introduces the preliminaries and problem formulation. Section IV presents the proposed framework. Simulation and hardware experiments are reported in Section V and Section VI, respectively. Concluding remarks are in Section VII.
@@ II Related Work
In this section, we will briefly review the most important related work on kinematic control of closed architecture and safe control. An overview of the comparison between our work and the most related works is presented in Table I.
A vast body of research on adaptive [slotine1987adaptive] and robust [slotine1985robust] torque control has been developed over the past several decades to address various uncertainties in robotic manipulators. However, we do not extensively cover this line of work in our literature review due to the unique nature of our problem, which focuses on a robotic manipulator with a closed-architecture system. A direct comparison between torque control and kinematic control is not fair, as torque control—implemented in the inner-loop layer of a cascaded control structure—is inherently more responsive than outer-loop approaches. Instead, we aim for our detailed formulation and proposed solution to motivate the adaptation and redesign of torque control methodologies to better accommodate our specific context. Furthermore, the challenge of dealing with unknown inner-loop controllers remains a promising direction for future research across all existing approaches, including both ours and adaptive control methods.
II-A Adaptive Control
Learning-based Adaptive Control: [ahanda2022adaptive, ahanda2022task, khan2024control] use adaptive outer-loop controller based on approximated kinematics, dynamics, and inner-loop controller using a neural network. As shown in Table I, these methods make the weakest assumption regarding the inner-loop controller, the same as our work: the $\Psi(\cdot)$ term represents an unknown structure that may include proportional, integral, and state-dependent compensation components. However, these approaches involve unexplainable learning processes (e.g., a large number of tuning parameters and troubleshooting difficulties).
Conventional Adaptive Control:[wang2019dynamic] proposes dynamic modularity, which employs an adaptive control method with a moderate assumption (see Table I): a known structure of the inner-loop controller, either PI or PID.
Note that these works focus only on tracking control tasks without considering safety. In our experiment, we will show a comparison of tracking performance with the learning-based adaptive control, since we share the same assumption of the inner-loop controller.
@@ II-A Adaptive Control
[ahanda2022adaptive, ahanda2022task, khan2024control]
@@
[molnar2021model, singletary2021safety, molnar2023safety]
Robust CBF - - [mrdjan2018robust, kolathaya2019input, alan2021safe, buch2022robust, nguyen2022robust, alan2023parameterized, wang2023disturbance, dacs2022robust, alan2023disturbance, zhao2020adaptive, sun2024safety, das2025robust, wang2024safety, cao2024safety, zhang2024eso, wang2023composite]
@@ II-B Reduced-order model-based CBF
To ensure the safety of robotic systems, many efforts using CBF theory focus on high-fidelity dynamic models with torque control. However, obtaining an accurate high-fidelity model is a fundamental challenge. Reduced-order model-based CBFs have been proposed to provide safety guarantees with a kinematic controller, even when using a low-fidelity model. [molnar2021model] aims to ensure safety using safe velocity without relying on a high-fidelity dynamical model. However, their safety guarantee is based on a strong assumption of a perfect inner-loop controller: any outer-loop velocity command can be tracked quickly. [singletary2021safety, molnar2023safety] propose kinetic energy-based CBFs to ensure safety at the kinematic control level, allowing for some robustness to model uncertainty, but still requiring many dynamic model parameters and a known inner-loop controller with a specific PD structure.
The fundamental difference between this line of research and our work is that we focus on a closed architecture with minimal available information about the inner-loop controller while allowing the dynamic model to exhibit significant uncertainty. As shown in Table I, due to the different assumptions, it is difficult to compare the safety performance of these works with ours.
@@ II-C Robust CBF
The original CBF [ames2016control] relies on a precise dynamic model to ensure that the system’s trajectory remains within a forward-invariant safety set, guaranteeing safety. To address this fundamental limitation in the presence of model uncertainty, robust CBFs have been proposed in recent years. These methods recover formal safety guarantees by incorporating additional terms in the constraints of CBF formulation to compensate for model uncertainties and external disturbances [mrdjan2018robust, kolathaya2019input, alan2021safe, buch2022robust, nguyen2022robust, alan2023parameterized, wang2023disturbance, alan2023disturbance, sun2024safety, das2025robust, wang2024safety, chen2023robust, zhao2020adaptive, dacs2022robust, zhang2024eso, wang2023composite, cao2024safety]. Among these methods, disturbance observer-based robust CBF (DOB-CBF) [chen2023robust, zhao2020adaptive, dacs2022robust, zhang2024eso, wang2023composite, cao2024safety, wang2023disturbance, alan2023disturbance, sun2024safety, das2025robust, wang2024safety] is one of the most popular approaches for handling uncertainties in dynamic models, as it allows for active disturbance estimation without requiring the worst-case disturbance bound [sariyildiz2020disturbance, sadeghian2013task]. As a result, DOB-CBF approaches, such as [wang2023disturbance], often exhibit less conservatism compared to the robust CBF methods that rely on worst-case disturbance assumptions [nguyen2022robust].
In [dacs2022robust], a DOB is designed to estimate the disturbance within the CBF dynamics, rather than the original system’s disturbance, to guarantee safety. To further enhance robustness and reduce conservativeness, [alan2023disturbance] integrates the DOB from [dacs2022robust] with the concept of tunable input-to-state safe control. Instead of using a disturbance observer to estimate disturbances within the CBF dynamics, [zhao2020adaptive] introduces an adaptive law to estimate disturbances in the original system dynamics. A sampled-data safe control strategy combined with a DOB is proposed in [sun2024safety] to address uncertainties caused by both intersample time intervals and disturbances. In [das2025robust], the original high-order CBF-QP is reformulated as a second-order cone program for systems with a specific structure, where the relative degree differences between the control inputs and disturbances are all less than or equal to 1. To handle unmatched disturbances in general nonlinear systems with disturbances, [wang2024safety] employs a DOB to estimate disturbances and the bounds of their higher-order derivatives, which are challenging to obtain in practice. In [cao2024safety], a novel safety-critical generalized predictive control method is introduced for speed regulation of a permanent magnet synchronous motor with current and voltage constraints, utilizing fixed-time sliding-mode disturbance observers to estimate disturbances. These approaches assume access to all true states. However, in practice, only measurements are often available. In such cases, state uncertainty needs to be accounted for, e.g., by incorporating an ESO within the CBF framework, as seen in [chen2023robust, zhang2024eso, wang2023composite].
Note that in these previous works, the estimation error bounds of disturbances are often relaxed using Lyapunov methods, whereas in our approach, they are obtained directly from the solution of the observer’s error dynamics. As a result, our method tends to be less conservative. We will provide more experimental comparisons on the tightness of the error bounds with state-of-the-art methods, such as [dacs2022robust].
@@ III Preliminary and Problem Formulation
In this section, we will briefly go through the necessary preliminary, then introduce our problem formulation.
@@ III-A System Dynamics
The most common inner-loop controller for robot manipulators typically consists of a PID controller or PD controller, often combined with dynamic compensation, such as gravity [kelly2005control]. For stability, most inner-loop controllers include a derivative term, $(\dot{q}-\dot{q}_{d})$, where $\dot{q}_{d}$ is a joint velocity command.
@@
with $K_{d}$ a proportional control gain, and $\Psi(\cdot)$ is a general, unknown function that may include proportional, integral, and dynamic compensation terms. To the best of our knowledge, this assumption is the most general in the literature (see Table I), requiring minimal knowledge of the inner-loop controller, similar to [ahanda2022adaptive], and more general than [wang2019dynamic].
@@
This assumption is mild in practice, as the geometric parameters of a robotic arm are easy to measure.
Then, the system dynamic model combining (1) and (2) can be expressed as:
$\begin{split}\ddot{q}&={M}^{-1}\Bigl(-{C}\dot{q}-{G}-F_{r}+{J}^{T}F_{ext}\\ &-{K}_{d}\dot{q}+\Psi(\cdot)+{K}_{d}\dot{q}_{d}\Bigr)\end{split}.$ (3)
Note that we have removed torque control, as it is not accessible in a closed architecture. $\dot{q}_{d}$ is the kinematic control that will be manipulated in the outer-loop layer, and its control law will be designed. Another reason for selecting $\dot{q}_{d}$ as the control input, rather than $q_{d}$, is to avoid the need for differentiation, which could introduce significant noise in practical applications.
@@ III-B Robot Kinematic Control with Known Model
If the friction term, $F_{r}$, and the external disturbance, $F_{ext}$, are neglected in (1), the exact system dynamics can be leveraged to design a control law:
@@
where $\ddot{q}^{\star},\dot{q}^{\star},q^{\star}\in\mathbb{R}^{n}$ represent the desired acceleration, velocity, and position of a reference joint trajectory generated by an upstream planner. $k_{p}\in\mathbb{R}^{n\times n}$ and $k_{d}\in\mathbb{R}^{n\times n}$ are the proportional and derivative gain matrices, respectively. Note that $\ddot{q}^{\star}$ serves only as a feedforward term, while $u_{0}$ remains a position controller.
By substituting (4) and (5) into (3), we get:
@@
By properly selecting $k_{p}$ and $k_{d}$, q{q} will converge to ${q}^{\star}$ asymptotically.
@@ III-C Control Barrier Functions
The robotic system (3) can be rewritten as a nonlinear control-affine system when ignoring the disturbance vector $F_{ext}$:
@@
The safety of system (7) can be guaranteed using the concept of a safety set. A set $\mathcal{C}$ is considered to be a safety set if it is forward invariant in the state space $\mathcal{X}$, i.e., for system (7) if solutions for some $u\in\mathcal{U}$ starting at any initial safe condition $x(0)\in\mathcal{C}$ satisfy $x(t)\in\mathcal{C}$, $\forall t\geq 0$. The safety set $\mathcal{C}$ is defined as a 0-superlevel set of a continuously-differential function $h(x):\mathbb{R}^{2n}\rightarrow\mathbb{R}$ as:
@@
(CBF [ames2016control]) A continuous and differentiable function $h:\mathbb{R}^{2n}\rightarrow\mathbb{R}$ is called a CBF for the system (7) with relative degree 1, if
$\sup\limits_{u\in\mathcal{U}}\left(L_{\psi}h(x)+L_{g}h(x)u\right)\geq-\beta(h(x))$ (9)
for all $t\geq 0$ and $x\in\mathcal{X}$, where $L_{\psi}h(x)\triangleq\frac{\partial h(x)}{\partial x}\psi(x)$, $L_{g}h(x)\triangleq\frac{\partial h(x)}{\partial x}g(x)$, and $\beta(s)$ is an extended class $\mathcal{K}$ function [^1] A continuous function: $\beta:[-b,a)\rightarrow[-\infty,\infty)$ for some $a>0,b>0$ is said to belong to extended class $\mathcal{K}$, if it is strictly increasing and $\beta(0)=0$..
The set of control signals ensuring the system safety is defined as:
$\begin{array}[]{r@{}l}K_{\text{cbf}}(t,x)\triangleq\{&u\in\mathcal{U}:L_{\psi}h(x)\\ &+L_{g}h(x)u\geq-\beta(h(x))\}.\end{array}$ (10)
For a system with an arbitrary relative degree, we need the following higher-order CBF.
Definition 2
(Exponential CBF (ECBF) [nguyen2016exponential, xiao2019control]) Consider system (7) with relative degree $r$ for an $r$-times continuously differentiable function $h$, i.e., $L_{g}L_{\psi}h(x)=\cdots=L_{g}L_{\psi}^{r-2}h(x)=0$ and $L_{g}L_{\psi}^{r-1}h(x)\neq 0,\forall x\in\mathcal{C}$. $h(x)$ is an ECBF if there exists a row vector $K_{a}\in\mathbb{R}^{r}$ satisfying $\forall x\in\mathcal{C}$
@@
where $\eta_{b}(x)=[h(x),\dot{h}(x),\cdots,{h^{(r-1)}(x)}]^{T}$, $K_{a}=[k_{1},\cdots,k_{r}]$, and the values of $k_{1},\cdots,k_{r}$ satisfy specific properties given in [nguyen2016exponential, xiao2019control].
In order to leverage a CBF to guarantee safety, the control problem is formulated as a quadratic program (QP) with a CBF as a hard constraint. The QP formulation is as follows [ames2016control]:
@@
Research Objective: Given the system described in (3), we have the following nominal dynamic system with minimal knowledge of the system and the inner-loop controller:
$\ddot{q}=\bar{M}^{-1}\Bigl(-\bar{C}\dot{q}-\bar{G}+\bar{K}_{d}\dot{q}_{d}\Bigr),$ (13)
where $\bar{M}$, $\bar{C}$, and $\bar{G}$ are the nominal model parameters, with $\bar{K}_{d}$ representing the nominal gain of the inner-loop controller.
The objective is to develop a controller for $\dot{q}_{d}$ that enables the system to accurately follow the reference trajectory and guarantee safety, such as collision avoidance in task space. This control design must address significant challenges, including: (i) an imperfect and largely unknown inner-loop controller, (ii) uncertain system dynamics, and (iii) the presence of external disturbances.
@@
We use the nominal model (14) for all the simulations and experiments in this paper.